# Phase transition and universality of the majority-rule model on complex networks

### 21 Feb 2024 v1

@@ 4 Final Remark
We conducted an analysis of the order-disorder phase transition and the universality class of the majority-rule model implemented on heterogeneous networks, including the Barabasi-Albert, Watts–Strogatz, and Erdos-Renyi networks, wherein each node maintains connections to at least two other nodes. Each agent possesses two possible opinions, namely ±$\pm 1$, randomly distributed across all network nodes. Additionally, each agent adheres to the principle of conformity, characterized by behavior aligned with the majority opinion. Notably, agents also exhibit anticonformity behavior (anticonformist agents) and independence behavior (independent agents). Within this model, anticonformist agents adopt the minority opinion, adjusting their opinion or state to align with the minority viewpoint, while independent agents act autonomously, unaffected by the influence of other agents. In a given population, three agents are selected randomly, and with a probability $p$, agents exhibit anticonformity and independence behaviors. For independent agents, the three agents change their opinions without group influence. In other situations, when all three agents hold identical opinions, they collectively alter their opinions, characteristic of anticonformist agents. Conversely, with a probability of $1-p$, the agents conform by adopting the majority opinion. In our investigation of the phase transition, we analyzed to estimate critical exponent parameters using finite-size scaling relations to ascertain the universality class associated with the model. It was observed that the critical point for the anticonformity model is smaller in magnitude compared to that of the independence model across all network configurations (see Table 1). Remarkably, for both anticonformity and independence models applied to all network topologies, the critical exponents exhibited striking similarities with those found in the opinion dynamics model defined on a complete graph and the mean-field Ising model. These results suggest that the models fall within the same universality class as the opinion dynamics model defined on a complete graph and the mean-field Ising model.

### 26 Feb 2024 v2

@@ 4 Final Remark
We conducted an analysis of the order-disorder phase transition and the universality class of the majority-rule model implemented on heterogeneous networks, including the Barabasi-Albert, Watts–Strogatz, and Erdos-Renyi networks, wherein each node maintains connections to at least two other nodes. Each agent possesses two possible opinions, namely ±$\pm 1$, randomly distributed across all network nodes. Additionally, each agent adheres to the principle of conformity, characterized by behavior aligned with the majority opinion. Notably, agents also exhibit anticonformity behavior (anticonformist agents) and independence behavior (independent agents). Within this model, anticonformist agents adopt the minority opinion, adjusting their opinion or state to align with the minority viewpoint, while independent agents act autonomously, unaffected by the influence of other agents.
In a given population, three agents are selected randomly, and with a probability $p$, agents exhibit anticonformity and independence behaviors. For independent agents, the three agents change their opinions without group influence. In other situations, when all three agents hold identical opinions, they collectively alter their opinions, characteristic of anticonformist agents. Conversely, with a probability of $1-p$, the agents conform by adopting the majority opinion.
In our investigation of the phase transition, we analyzed to estimate critical exponent parameters using finite-size scaling relations to ascertain the universality class associated with the model. It was observed that the critical point for the anticonformity model is smaller in magnitude compared to that of the independence model across all network configurations (see Table 1). Remarkably, for both anticonformity and independence models applied to all network topologies, the critical exponents exhibited striking similarities with those found in the opinion dynamics model defined on a complete graph and the mean-field Ising model. These results suggest that the models fall within the same universality class as the opinion dynamics model defined on a complete graph and the mean-field Ising model.